OpenStudy (rowanmartinfan):

who knows how to get the absolute value of complex numbers if so plz tell me

OpenStudy (anonymous):

\[\left| A+Bi \right|=\sqrt{A^2+B^2}\]

OpenStudy (rowanmartinfan):

yea i just found that out lol

OpenStudy (anonymous):

Okay (:

OpenStudy (rowanmartinfan):

brb

OpenStudy (anonymous):

ok

OpenStudy (rowanmartinfan):

ok im back again